OpenStudy (anonymous):

If csc alpha = 7/3 and cot alpha= 2 square root of 10 over 3 find sec alpha

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

think of cosecant as "hypotenuse over adjacent" and draw a triangle

OpenStudy (misty1212):

find the missing side via pythagoras then you can take any trig ratio you like

OpenStudy (anonymous):

After I find the missing side do I multiply them and use the keep change flip method?

OpenStudy (misty1212):

not sure what you are asking

OpenStudy (misty1212):

other side is \[\sqrt{7^2-3^2}=\sqrt{49-9}=\sqrt{40}=2\sqrt{10}\]

OpenStudy (anonymous):

For the missing side it was square root of 40

OpenStudy (misty1212):

yeah \(\sqrt{40}\) or \(2\sqrt{10}\)

OpenStudy (misty1212):

oh wow i made a huge mistake

OpenStudy (anonymous):

What happens after that how do I find sin if both cot and csc is 2 square root of 10

OpenStudy (misty1212):

cosecant is the reciprocal of sine not cosine

OpenStudy (anonymous):

Yea

OpenStudy (misty1212):



OpenStudy (misty1212):

secant is the reciprocal of cosine, so it is hypotenuse over adjacent i.e. \[\frac{7}{2\sqrt{10}}\]

OpenStudy (anonymous):

So we multiply square root of ten on both top and bottom to cancel the square root and it gives me the final answer of 7 square root 10 over 20 thank u so much!!!!